Raghvan has spent a quarter of his life as a boy.One fifth as a youth,one-third as a man and 13 in old age.what is his present age?
Answer by mananth(16946)

Raghvan has spent a quarter of his life as a boy.One fifth as a youth,one-third as a man and 13 in old age.what is his present age?
let his present age be x
1/4 as boy
1/5 as youth
1/3 as man
13 in old age
1/4 *x + 1/3 * x+ 1/5* x +13 =x
1/4 *x + 1/3 * x+ 1/5* x -x = -13
take LCM
15x+20x+12x-60x / 60 = -13
15x+20x+12x-60x=-13* 60
-13x=-780
x=60
Raghavan's age is 60 years
Ananth